Fermi National Laboratory


Accelerator Update for November 9, 2005 - November 11, 2005

Wednesday November 9

The day shift began with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) in access, with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Pbar experts began a study period at 9:19 AM.

CDF came out of access at 9:50 AM. D0 came out of access at 10:38 AM.

Operations put the Tevatron (TeV) through a dry squeeze at 11:21 AM.

Operations put the TeV into shot setup at 11:56 AM.

NuMI experts reported at 12:01 PM that their orbit verification program was now working. Operators unmasked two position monitors HP608N and VP607N and then NuMI began tripping off due to bad positions. Another expert arrived to help investigate. Operators re-masked HP608N and VP607N.

The Pelletron cooling work was completed at 1:44 PM.

At 1:52 PM, Operations reported that the NuMI orbit verifier failed again. An expert investigated.

Operations established Recycler store 4494 at 1:55 PM with an initial luminosity of 112.89E30.

Operations began preparing the Recycler to stash at 2:18 PM.

A NuMI expert reported that he had fixed the orbit verifier at 2:47 PM.

The Recycler stash ended. Pbar, MTest, and MCenter resumed normal operations at 3:15 PM.

After a NuMI expert tested the orbit verifier, Operations resumed sending beam to NuMI at 3:45 PM.

A Main Injector (MI) expert began a 2.5 Mhz study around 4 PM.

At 4:13 PM, Operations reported that the NuMI target profile monitor had stopped operating. Operations contacted an instrumentation expert.

Operations reported at 5:11 PM that with the Pelletron working the Electron Cooling was coming back on line.

At 8:24 PM, a Booster kicker (MKS02) tripped off and halted Switchyard (SY) beam.

At 8:40 PM, a MI RF station (MIRF18) tripped off.

At 8:45 PM, a Booster RF station (BRF12) tripped off for the first of three times in forty-three minutes. It reset each time.

MI studies ended at 9:54 PM.

The Recycler began preparing to stash at 9:57 PM. The stash ended. Pbar, MTest, and MCenter resumed normal operations at 10:52 PM.

Thursday November 10

The midnight shift began with Operations monitoring store 4494 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

MTest experimenters reported at 1:41 AM that they were finished with beam for the night.

The Recycler began preparing to stash at 2:51 AM.

At 2:55 AM, Operations reported a P1 transfer line vacuum burst.

The stash ended. Pbar, MTest, and MCenter resumed normal operations by 3:54 AM.

The Recycler began preparing to stash at 9:29 AM.

Pbar experts decided to use this stash as a study period for the P1, P2, AP1, and AP3 transfer lines.

The Pbar study ended at 12:32 PM. The Recycler stash resumed.

The stash ended. Pbar, MTest, and MCenter resumed normal operations by 1:12 PM.

TeV experts began an end-of-store study at 2:12 PM.

MI studies ended at 4:26 PM.

At 4:55 PM, the P1 line suffered another vacuum burst. Operators managed to stop its cascade by closing valves. The vacuum pump was reset and turned back on.

Operations terminated store 4494 at 5:18 PM.

Operations put the TeV into shot setup at 5:52 PM.

Operations established Recycler store 4495 at 8:12 PM with an initial luminosity of 166.8E30. This store set a New Record Luminosity. Pbar, MTest, and MCenter resumed normal operations by 8:20 PM.

At 8:45 PM, Pbar and SY beam halted due to tripped load transfer switch (H703) for the A1 transfer line. Operators investigated and then called in an expert. He had the switch working by 10:20 PM.

MCenter resumed taking beam at 10:28 PM.

The Recycler began preparing to stash at 11:17 PM.

Friday November 11

The midnight shift began with Operations monitoring store 4495, with the Recycler was stashing, with Pbar and SY off due to stashing, and with MiniBooNE and NuMI taking beam.

At 12:14 AM, a Linac RF station (LRF3) tripped off. Operators managed to get it back on line fifteen minutes later.

The stash ended. Pbar and SY resumed normal operations at 12:24 AM.

At 12:28 AM, LRF3 tripped off for the first of five times in an hour. Operators managed to reset it each time.

At 2:17 AM, LRF3 tripped off and wouldn’t reset. Operators called in an expert, but then managed to reset the station at 2:32 AM.

At 5:04 AM, LRF3 tripped off twice. It wouldn’t reset on the second trip. Operators responded along with the expert. It then tripped off four more times over the next ten minutes.

The Recycler began preparing to stash at 5:27 AM.

The stash ended. Pbar, MTest, and MCenter resumed normal operations by 6:15 AM.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow LRF3 repairs this morning followed by a reverse proton study. MI will conduct a study during the reverse proton study.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/11/2005   email Fermilab